# PECCARY: A novel approach for characterizing orbital complexity, stochasticity, and regularity

### v1

@@ 3.1 Setting up PECCARY
To use PECCARY, the code can be installed from the Python Package Index (PyPI) via the command pip install peccary or downloaded from the GitHub repository. [^1] https://github.com/soleyhyman/chaos-orbits Documentation and tutorials for running the code can be found on the PECCARY website. [^2] https://peccary.readthedocs.io At its most basic, all that is needed is a time-series and a chosen sampling interval $\ell$. By default, the sampling size is set to $n=5$ (see Section 2.1).
@@ 3.3 Interpreting PECCARY Values
Chaotic systems, on the other hand, have a characteristic shape to their curves that depend on whether they are discrete or continuous in nature. Discrete chaotic systems have a characteristic timescale that is inherently set to be $\ell=1$. As such, the maximum value for the Statistical Complexity occurs when $\ell=1$. By contrast, the maximum Statistical Complexity for continuous chaotic system depends on the approximate natural timescale. In terms of $H$- and $C$-curves, the value for Permutation Entropy increases with increasing sampling interval, while the value for Statistical Complexity increases to some maximum value at a particular pattern timescale, and then decreases. Examples of both discrete and chaotic maps are given in Sections 4.1.3 and 4.2.
@@ 4.1.2 Noise varieties
Figure 7: Permutation Entropy and Statistical Complexity values for a range of sampling intervals for five different colors of noise (pink, red, violet, blue, and white). Left panel: Permutation Entropy as a function of sampling interval/pattern timescale. All $H$ values are very high (close to $H=1$), indicating the presence of close to all possible permutations of patterns. Horizontal dashed lines represent the upper and lower limits of $H_{\rm per}^{\rm min}(n=5)$ and $H_{\rm per}^{\rm max}(n=5)$, respectively. Center panel: Statistical Complexity as a function of sampling interval/pattern timescale. $C$ values are very low for all colors of noise, indicating that the pattern probability distributions are close to uniform. Right panel: The $[H,C]$ values for sampling interval of $\ell=1$ plotted on the $HC$-plane for all noise varieties fall well within the stochastic region.
@@ 4.1.3 Chaotic systems
Two well-studied examples of chaos are the Hénon map (Eq. 11) and the Lorenz strange attractor. The Hénon map is a discrete chaotic map, while the Lorenz strange attractor is continuous. The Lorenz strange attractor is a system described by (Lorenz, 1963),
@@
Alternately, the timescale at which the maximum Statistical Complexity ($C$) occurs can be used for the ideal sampling. In this method, the sampling interval corresponding to the peak value in the $C(\ell)$ curve is determined. From that value, the pattern timescale $t_{\rm pat}$ can be calculated (via Equation 1). Depending on the $t_{\rm pat}/t_{\rm nat}$ ratio used (typically 0.4), that $t_{\rm pat}$ value can be used to find the natural oscillatory timescale $t_{\rm nat}$. This is called the “$t_{\rm nat}$ from maximum $C(\ell)$” method.
@@ 4.3 Astrophysical Examples
The tracer particle simulations for the astrophysical examples in the following subsections were created with galpy (Bovy, 2015), using the symplec4_c integrator.
@@ 4.3.1 Keplerian Potential
While there are many types of regular orbits in astrophysics, only the point-mass (i.e., Keplerian) potential produces a special case of non-circular orbits that close in a single period, due to the fact that the radial and azimuthal frequencies are equal. As such, the Keplerian potential is an ideal scenario for testing regular orbits that close after $2\pi$. In this case, galpy’s Orbit.from_name(’solar system’) function was used to create a tracer-particle simulation of the orbits of the eight planets of the solar system for a duration of 100 years, at a resolution of 2.85 days. Figure 11 demonstrates that all the values fall on and within the periodic/regular boundary of the $HC$-plane when using the x-coordinate for the orbits.
Figure 11: $HC$-plane showing the $[H,C]$ coordinates for tracer particle simulation x-coordinates of the Solar System when $t_{\rm pat}/t_{\rm nat}=0.4$. All data points fall within the periodic/regular boundary, consistent with the fact that these are circular orbits.
@@ 4.3.2 Globular Cluster
A spherical potential is a minimally intricate astrophysical example for testing the PECCARY method. Using a spherical Plummer potential (galpy.potential.PlummerPotential) and self-consistent isotropic and spherical Plummer distribution function (galpy.df.isotropicPlummerdf), $10^{4}$ tracer particles (representing stars) were evolved for a duration of 1 Gyr and an orbit integration time resolution of 0.1 Myr. Figure 12 shows a sampling of 50 stellar orbits plotted on the $HC$-plane based on the $[H,C]$ values calculated from the x-coordinates of orbits with $t_{\rm dur}/t_{\rm pat}\geq 1.5$ and $t_{\rm pat}/t_{\rm nat}=0.4$. Many orbits fall on or within the periodic/regular boundary, with others falling within the “regular” regime. The divergence from the periodic boundary line is likely due to the fact that these are not closed orbits. The PECCARY method will be developed further for investigating regimes such as this in a future paper.
Figure 12: $HC$-plane showing the $[H,C]$ values when $t_{\rm pat}/t_{\rm nat}=0.4$ for globular cluster tracer particles. The x-coordinates are used for the PECCARY method, with orbital periods satisfying the $t_{\rm dur}/t_{\rm nat}\geq 1.5$ requirement. Most of the $[H,C]$ values fall on or within the periodic/regular boundary/region, though some have lower complexity values. None of the $[H,C]$ values for the orbits exceed the $H_{\rm per}^{\rm max}$ limit for $n=5$, consistent with the expectation that all of the orbits for a spherical potential are regular.
@@ 4.3.3 Triaxial Halo
A triaxial potential will exhibit regular orbits, as well as chaotic orbits. To verify that PECCARY returns this same conclusion, a tracer-particle simulation of a Navarro-Frenk-White (NFW) potential (galpy.potential.TriaxialNFWPotential) was created with an isotropic NFW distribution function that samples a spherical NFW halo (galpy.df.isotropicNFWdf). The default galpy settings were used for the triaxial prescription, with the exception of the y/x and z/x axis ratios. These two ratios are referred to as $b$ and $c$, respectively, and were set to $b=1.66$ and $c=3$ with a normalization of 0.35 based on the reported values for a Milky Way-like halo in Rojas-Niño et al. (2012); Hesp & Helmi (2018). $10^{4}$ stars were evolved for a duration of 100 Gyr at a time resolution of 15.625 Myr.
Figure 13 shows $[H,C]$ values at $t_{\rm pat}/t_{\rm nat}=0.4$ for the x-coordinates of simulated stars with $t_{\rm dur}/t_{\rm nat}\geq 1.5$. Several of the orbits fall within the regular regime, where many of these were visually identified as tube orbits. Some orbits, particularly those with shorter orbital periods, have time-series characteristics that are highly complex. Future work will compare the results from PECCARY to other known diagnostics of chaos.
@@
Figure 13: $HC$-plane showing the $[H,C]$ values when $t_{\rm pat}/t_{\rm nat}=0.4$ for triaxial NFW halo tracer particles. The x-coordinates are used for the PECCARY method, with orbital periods satisfying the $t_{\rm dur}/t_{\rm nat}\geq 1.5$ requirement. As expected for a triaxial potential, there is a mix particles falling in both the regular and complex zones.
4.3.4 Injected Noise
To test the sensitivity of PECCARY on several idealized examples of “noisy data,” different levels of noise were injected to the various orbits presented in Section 4.3.1-4.3.3. The noise injections were determined by calculating the approximate amplitude of each orbital x-coordinate time-series (i.e., maximum absolute value less the average value) and then adding a random value to each time-step. The random “errors” were drawn from a uniform distribution ranging from -1 to 1 and were then scaled by the amplitude of the time-series and the fraction of the noise (i.e., 0.1, 0.5, or 1). The noise injection “percentages” are defined as the amplitude of the noise relative to the amplitude of the signal.
Figure 14 compares values on the $HC$-planes of the three noise levels at sampling intervals equivalent to $t_{\rm pat}/t_{\rm nat}=0.4$ and for orbits with $t_{\rm dur}/t_{\rm nat}\geq 1.5$. Compared to the original datasets, periodic and regular orbits with injected noise appear to increase primarily in $H$, while complex orbits appear to have similar $[H,C]$ coordinates. To the level of 10% noise, the $[H,C]$ values have shifted significantly compared to the pure signals, although they do not appear as stochastic. At 50% and 100% noise, almost all or all of the $[H,C]$ points fall primarily within the stochastic region.
Figure 14: $HC$-planes showing the effect of noise injection on the PECCARY method for the three astrophysical example potentials. Circle, square, and triangle markers represent the Solar System, globular cluster, and triaxial halo simulations respectively. Left panel: Original $[H,C]$ values, as shown in Figures 11-13. Left-center panel: $[H,C]$ values for the three potentials with noise amplitude at 10% of the signal amplitude. Regular orbits appear to fall either within the regular region or below the primary complex region. Complex orbits still fall within the region of maximum complexity. Right-center panel: $[H,C]$ values for the three potentials with noise amplitude at 50% of the signal amplitude. Almost all of the values appear to fall within the stochastic region. Right panel: $[H,C]$ values for the three potentials with noise amplitude equal to the signal amplitude. All of the values to fall within the stochastic region.
@@ 5 Future Work
This paper introduces the PECCARY method for usage in astrophysics. While the measures of Permutation Entropy and Statistical Complexity have been used in other fields, including plasma physics (Weck et al., 2015), to great success, those bodies of work involved systems or models that were inherently discrete and could use a sampling interval of $\ell=1$. The fundamentally continuous nature of many astrophysical systems, as well as the varied origin of stochasticity (i.e., natural noise or background sources/behaviors), require great care with choosing the appropriate sampling schemes. The PECCARY method provides the first clear recommendations for using Permutation Entropy and Statistical Complexity measures to characterize the behaviors of continuous systems.
This study investigates how well the method works for several astrophysical systems with known behaviors (Section 4.3), but additional work is needed for widespread uses. Future papers will develop a method for estimating the confidence of the periodic/regular/chaotic/stochastic diagnosis and run comparisons with existing methods for chaos identification, such as frequency analysis mapping (Laskar, 1990; Valluri & Merritt, 1998; Valluri et al., 2012, 2016; Beraldo e Silva et al., 2019, 2023) and potentially Lyapanov exponents to test the robustness of PECCARY in this regime.
While Section 4.3.4 demonstrated an example of PECCARY’s ability to handle noise injected into a known system, additional work is needed on understanding the sensitivities of the method for characterizing orbital behavior in a noisy signal. Future research will center on developing more intricate astrophysical simulations (e.g., evolving, time-dependent potentials and large-scale n-body simulations) to test the PECCARY method. Other work will explore the efficacy of stacking time-series in order to improve reliability for shorter duration simulations and windowing, for understanding how a system changes dynamically over time.
@@ 6 Conclusions
Extensive documentation and examples for the corresponding PECCARY Python package are available online.333 https://peccary.readthedocs.io The source code can be found on GitHub [^4] https://github.com/soleyhyman/peccary and builds can be found on the PyPI project page. [^5] https://pypi.org/project/peccary/
The astrophysical simulations in Section 4.3 were created using High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department. SOH and KJD would like to acknowledge support for the above resources. KJD also acknowledges support provided by the Heising Simons Foundation grant # 2022-3927.

### v3

@@ 3.1 Setting up PECCARY
To use PECCARY, the code can be installed from the Python Package Index (PyPI) via the command pip install peccary or downloaded from the GitHub repository (S. Hyman & D. Schaffner, 2024). [^1] https://github.com/soleyhyman/peccary Documentation and tutorials for running the code can be found on the PECCARY website. [^2] https://peccary.readthedocs.io At its most basic, all that is needed is a time-series and a chosen sampling interval $\ell$. By default, the sampling size is set to $n=5$ (see Section 2.1).
@@ 3.3 Interpreting PECCARY Values
Chaotic systems, on the other hand, have a characteristic shape to their curves that depend on whether they are discrete or continuous in nature. A discrete, recursive series that is chaotic has an inherent timescale that is an integer value of 1. As such, the maximum value for the Statistical Complexity occurs when $\ell=1$. By contrast, the maximum Statistical Complexity for continuous chaotic systems depends on the approximate natural timescale. In terms of $H$- and $C$-curves, the value for Permutation Entropy increases with increasing sampling interval, while the value for Statistical Complexity increases to some maximum value at a particular pattern timescale, and then decreases. Examples of both discrete and chaotic maps are given in Sections 4.1.3 and 4.2.
@@ 4.1.2 Noise varieties
Figure 7: Permutation Entropy and Statistical Complexity values for a range of sampling intervals for five different colors of noise (pink, red, violet, blue, and white). Left panel: Permutation Entropy as a function of sampling interval/pattern timescale. All $H$ values are very high (close to $H=1$), indicating the presence of close to all possible permutations of patterns. Horizontal dashed lines represent the upper and lower limits of $H_{\rm per}^{\rm min}(n=5)$ and $H_{\rm per}^{\rm max}(n=5)$, respectively. Center panel: Statistical Complexity as a function of sampling interval/pattern timescale. $C$ values are very low for all colors of noise, indicating that the pattern probability distributions are close to uniform. Right panel: The $[H,C]$ values for sampling interval of $\ell=1$ plotted on the $HC$-plane for all noise varieties fall well within the stochastic region. Note that Brownian noise has slightly lower Permutation Entropy and slightly higher Statistical Complexity than some of the other noise colors due to its steep power spectrum slope, which results in the suppression of higher frequencies and results in less overall scatter in the resulting time-series.
@@ 4.1.3 Chaotic systems
Two well-studied examples of chaos are the Hénon map (Eq. 11) and the Lorenz strange attractor. The Hénon map is a discrete chaotic map, while the Lorenz strange attractor is continuous. The Lorenz strange attractor is a system described by (E. N. Lorenz, 1963),
@@
Alternatively, the timescale at which the maximum Statistical Complexity ($C$) occurs can be used for the ideal sampling. The use of the maximum Statistical Complexity ensures that the classification of the overall behavior of the system is as accurate as possible, since poorly-sampled regular time-series will tend towards the linear regime and poorly-sampled chaotic time-series will tend towards the stochastic regime. In this method, the sampling interval corresponding to the peak value in the $C(\ell)$ curve is determined. From that value, the pattern timescale $t_{\rm pat}$ can be calculated (via Equation 1). Depending on the $t_{\rm pat}/t_{\rm nat}$ ratio used (typically 0.4), that $t_{\rm pat}$ value can be used to find the natural oscillatory timescale $t_{\rm nat}$. This is called the “$t_{\rm nat}$ from maximum $C(\ell)$” method.
@@ 4.3 Astrophysical Examples
The tracer particle simulations for the astrophysical examples in the following subsections were created with galpy (J. Bovy, 2015), using the symplec4_c integrator.
@@ 4.3.1 Keplerian Potential
While there are many types of regular orbits in astrophysics, only the point-mass (i.e., Keplerian) potential produces a special case of non-circular orbits that close in a single period, due to the fact that the radial and azimuthal frequencies are equal. As such, the Keplerian potential is an ideal scenario for testing regular orbits that close after $2\pi$. In this case, galpy’s Orbit.from_name(‘solar system’) function was used to create a tracer-particle simulation of the orbits of the eight planets of the solar system for a duration of 100 years, at a resolution of 2.85 days. Figure 11 demonstrates that all the values fall on and within the periodic/regular boundary of the $HC$-plane when using the radial coordinates for the orbits. The natural timescales used in the PECCARY calculations for Figure 11 are the radial periods of the planet orbits, determined with galpy’s Orbit.Tr function.
Figure 11: $HC$-plane showing the $[H,C]$ coordinates for tracer particle simulation radial coordinates of the Solar System when $t_{\rm pat}/t_{\rm nat}=0.4$. The natural timescales ($t_{\rm nat}$) are the radial periods of the orbits, determined with galpy’s Orbit.Tr function. All data points fall within the periodic/regular boundary, consistent with the fact that these are circular orbits.
@@ 4.3.2 Globular Cluster
A spherical potential is a minimally intricate astrophysical example for testing the PECCARY method. Using a spherical Plummer potential (galpy.potential.PlummerPotential) and self-consistent isotropic and spherical Plummer distribution function (galpy.df.isotropicPlummerdf), $10^{4}$ tracer particles (representing stars) were evolved for a duration of 1 Gyr and an orbit integration time resolution of 0.1 Myr. Figure 12 shows a sampling of 50 stellar orbits plotted on the $HC$-plane based on the $[H,C]$ values calculated from the radial coordinates of orbits with $t_{\rm dur}/t_{\rm pat}\geq 1.5$ and $t_{\rm pat}/t_{\rm nat}=0.4$. The natural timescales ($t_{\rm nat}$) used are the radial periods of the star orbits, determined with galpy’s Orbit.Tr function. As demonstrated in Figure 12, all orbits fall on the periodic/regular boundary of the $HC$-plane.
Figure 12: $HC$-plane showing the $[H,C]$ values when $t_{\rm pat}/t_{\rm nat}=0.4$ for a sampling of 50 globular cluster tracer particles. The natural timescales ($t_{\rm nat}$) are the radial periods of the orbits, determined with galpy’s Orbit.Tr function. The radial coordinates are used for the PECCARY method, with orbital periods satisfying the $t_{\rm dur}/t_{\rm nat}\geq 1.5$ requirement. Most of the $[H,C]$ values fall on or within the periodic/regular boundary/region, though some have lower complexity values. None of the $[H,C]$ values for the orbits exceed the $H_{\rm per}^{\rm max}$ limit for $n=5$, consistent with the expectation that all of the orbits for a spherical potential are regular.
@@ 4.3.3 Triaxial Halo
A triaxial potential will exhibit regular orbits, as well as chaotic orbits. To verify that PECCARY returns this same conclusion, a tracer-particle simulation of a toy Navarro-Frenk-White (NFW) potential (galpy.potential.TriaxialNFWPotential), was created with an isotropic NFW distribution function that samples a spherical NFW halo (galpy.df.isotropicNFWdf). The triaxial NFW potential is defined by galpy as
$\rho(x,y,z)=\frac{\mathcal{A}}{4\pi a^{3}}\frac{1}{(m/a)(1+m/a)^{2}}\;,$ (15)
where $\mathcal{A}$ is the amplitude and $a$ is the scale radius. The parameter $m$ is described by
$m^{2}=x^{2}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}\;,$ (16)
where $x$, $y$, and $z$ are the Cartesian coordinates.
The parameters $b$ and $c$ are the y-axis/x-axis and z-axis/x-axis ratios of the density, respectively. The default galpy settings were used for the triaxial prescription, such that the amplitude of the potential is $\mathcal{A}=1$ and the scale length is set to $a=2$. The default values for the axis ratios were changed to $b=0.7$ and $c=0.5$, which fall within the range of mean values of Milky-Way-like halos in the Illustris-Dark (dark-matter-only) simulation, as reported in K. T. E. Chua et al. (2019). $10^{4}$ tracer particles were evolved for 10,000 time-steps in default galpy natural time units.
Figure 13 shows the $[H,C]$ values at $t_{\rm pat}/t_{\rm nat}=0.4$ for the r-coordinates of a sampling of 1300 orbits with $t_{\rm dur}/t_{\rm nat}\geq 1.5$ (i.e., with radial oscillation periods up to $6$ galpy time-steps). Due to the non-axisymmetric nature of the potential, the galpy Orbit.Tr calculations were not as reliable. Instead, the natural timescales ($t_{\rm nat}$) were approximated by finding the best fit for a sine wave to each time-series. Orbits were classified as either regular or chaotic/complex, based on their $[H,C]$ values. There are both regular and chaotic orbits present, with the number of regular orbits dominating. Future work will compare the results from PECCARY to other known diagnostics of chaos.
@@
Figure 13: $HC$-plane showing the $[H,C]$ values when $t_{\rm pat}/t_{\rm nat}=0.4$ for for a sampling of 1300 triaxial NFW halo tracer particles. The natural timescales ($t_{\rm nat}$), which are given in galpy natural units, are the approximated radial periods of the orbits, which are calculated by fitting a sine wave to the time-series. The radial coordinates are used for the PECCARY method, with orbital periods satisfying the $t_{\rm dur}/t_{\rm nat}\geq 1.5$ requirement. As expected for a triaxial potential, there is a mix of particles falling in both the regular and complex zones.
@@ 5 Future Work
This paper introduces the PECCARY method for usage in astrophysics. While the measures of Permutation Entropy and Statistical Complexity have been used in other fields, including plasma physics (P. J. Weck et al., 2015), to great success, those bodies of work involved systems or models that were inherently discrete and could use a sampling interval of $\ell=1$. The fundamentally continuous nature of many astrophysical systems, as well as the varied origin of stochasticity (i.e., natural noise or background sources/behaviors), require great care with choosing the appropriate sampling schemes. The PECCARY method provides the first clear recommendations for using Permutation Entropy and Statistical Complexity measures to characterize the behaviors of continuous systems.
This study investigates how well the method works for several astrophysical systems with known behaviors (Section 4.3), but additional work is needed for widespread uses. Future papers will develop a method for estimating the confidence of the periodic/regular/chaotic/stochastic diagnosis and run comparisons with existing methods for chaos identification, such as frequency analysis mapping (J. Laskar, 1990; M. Valluri & D. Merritt, 1998; M. Valluri et al., 2012, 2016; L. Beraldo e Silva et al., 2019, 2023) and potentially Lyapanov exponents to test the robustness of PECCARY in this regime.
Additional work is also needed to understand the sensitivities of the method for characterizing orbital behavior in a noisy signal for both idealized and realistic systems/simulations. Future research in this area will test how PECCARY responds to injecting different levels of noise in frozen n-body simulations before expanding that to more intricate astrophysical simulations (e.g., evolving, time-dependent potentials and large-scale n-body simulations). Other work will explore the efficacy of stacking time-series in order to improve reliability for shorter duration simulations and windowing, for understanding how a system changes dynamically over time.
@@ 6 Conclusions
Extensive documentation and examples for the corresponding PECCARY Python package (S. Hyman & D. Schaffner, 2024) are available online. [^3] https://peccary.readthedocs.io The source code can be found on GitHub [^4] https://github.com/soleyhyman/peccary and builds can be found on the PyPI project page. [^5] https://pypi.org/project/peccary/
The authors thank Michael Petersen and Leandro Beraldo e Silva for their stimulating conversations and suggestions. SÓH would like to acknowledge support from the University of Arizona’s Theoretical Astrophysics Program’s Travel Grant. KJD acknowledges support provided by the Heising-Simons Foundation grant # 2022-3927. The astrophysical simulations in Section 4.3 were created using High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department. We respectfully acknowledge the University of Arizona is on the land and territories of Indigenous peoples. Today, Arizona is home to 22 federally recognized tribes, with Tucson being home to the O’odham and the Yaqui. Committed to diversity and inclusion, the University strives to build sustainable relationships with sovereign Native Nations and Indigenous communities through education offerings, partnerships, and community service. We recognize the Lenape Indian tribe as the original inhabitants of eastern Pennsylvania, where Bryn Mawr College stands. We acknowledge the Lenape people as the indigenous stewards of their homelands and also the spiritual keepers of the Lenape Sippu, or Delaware River. We respect and honor the ancestral caretakers of the land, from time immemorial until now, and into the future.